OpenStudy (anonymous):

what is the midpoint, of line BC if... B is (11, 20) and C is (0, 20). Please help!

terenzreignz (terenzreignz):

Hello :) To get the midpoint of the line SEGMENT BC (Lines don't have midpoints, they have infinite length!) Where \[\large B=(x_1,y_1) \ \ \ C=(x_2,y_2)\] You just get the average of their x and y coordinates... Midpoint is given by \[\large M=(\frac{x_1+x_2}{2},\frac{y_1+y_2}{2})\]

OpenStudy (anonymous):

Then it's (5, 20). Correct?

terenzreignz (terenzreignz):

How do you get 5?

OpenStudy (anonymous):

0+ 10=10 10/2= 5

terenzreignz (terenzreignz):

why 10? It's 11. That's what you posted, at least (11 , 20) ( 0 , 20)